OpenStudy (anonymous):

Write the equation in slope-intercept form. What are the slope and y-intercept? -2x-11y=5

OpenStudy (anonymous):

y = -2/11x - 5/11 so, slope m = -2/11 y-intercept is -5/11

OpenStudy (anonymous):

Thank you for your help!